# Toward Relative Positional Encoding in Spiking Transformers

### v2

@@ Appendix C Performance of QKFormers on Image Classification
Table S2: Accuracy (%) of QKFormer on image classification benchmarks. Numbers with denote our implementations. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. Results are averaged across $4$ random seeds.
@@ D.2 Time-series Forecasting
$\mathrm{RSE}$ $=\sqrt{\frac{\sum_{m=1}^{M}||\mathbf{Y}^{m}-\hat{\mathbf{Y}}^{m}||^{2}}{\sum_{m=1}^{M}||\mathbf{Y}^{m}-\bar{\mathbf{Y}}||^{2}}}.$ (25)
@@
All SNNs take $4$ time steps for spiking neurons. We construct all Spikformer as $2$ blocks, setting the feature dimension as $256$, and the hidden feature dimension in FFN as $1024$. As for QKFormer, we set the block number as $4$, $2$ of which are QK blocks and the other $2$ are Spikformer blocks.
@@ D.4 Image Classification
Appendix E Limitations and Future Work
E.1 Limitations
@@ 8. Answer: [Yes]
  9. Justification: We have discussed the limitations and future work in Appendix E.
@@ 33. Answer: [Yes]
  34. Justification: Our reported results are all averaged over several random seeds. We have reported the standard deviation of the results in Table 2.
@@ 58. Answer: [Yes]
  59. Justification: The datasets we used in the paper are all public datasets. Please refer to Appendix D for details of datasets.

### v3

@@ Appendix C Performance of QKFormers on Image Classification
Table S2: Accuracy (%) of QKFormer on image classification benchmarks. Numbers with denote our implementations. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. Results are averaged across 44 random seeds.
@@ D.2 Time-series Forecasting
[…]
@@
All SNNs take $4$ time steps for spiking neurons. We construct all Spikformer as $2$ blocks, setting the feature dimension as $256$, and the hidden feature dimension in FFN as $1024$. As for QKFormer, we set the block number as $4$, 22 of which are QK blocks and the other $2$ are Spikformer blocks.
@@ D.4 Image Classification
Appendix E Analysis on Hardware-Friendliness of Log-PE
First, Log-PE doesn’t need to perform logarithmic operations directly on hardware during inference. Specifically, the relative position bias is defined as $\mathbf{R}_{i,j}=\begin{bmatrix}\lceil\log_{2}\left(\frac{L-1}{|i-j|+1}\right)\rceil\end{bmatrix}$, where $\lceil\cdot\rceil$ denotes the ceiling (round-up) function, and $L$ is the maximum sequence length. Since this bias depends only on the relative positions and the predefined sequence length, the entire bias matrix can be computed offline and stored ahead of time, eliminating the need for any runtime computation.
Secondly, even if one wishes to compute the logarithmic transformation on hardware, this can be efficiently achieved using a lookup table (LUT) implementation. Given an unsigned integer input of $N$ bits, we partition the input range into $K$ intervals. Each interval is approximated using a piecewise linear function $y=ax+b$, with the parameters $(a,b)$ stored in the LUT. The total LUT storage cost is: $K\cdot(N+2P)\ \text{bits}\approx\frac{K\cdot(N+2P)}{8}\ \text{bytes},$ where $P$ is the bit width of the parameters.
This strategy is similar to existing SNN approximations for exponential/leaky functions and has been successfully deployed in many types of neuromorphic chips, such as Intel Loihi [35]. Hence, the hardware implementation of Log-PE is efficient, low-cost, and practically feasible.
Appendix F Limitations and Future Work
F.1 Limitations
@@ 8. Answer: [Yes]
  9. Justification: We have discussed the limitations and future work in Appendix F.
@@ 33. Answer: [Yes]
  34. Justification: Our reported results are all averaged over several random seeds. We have reported the standard deviation of the results in Table˜2.
@@ 58. Answer: [Yes]
  59. Justification: The datasets we used in the paper are all public datasets. Please refer to Appendix˜D for details of datasets.